Determine the size of spherical balloon required to lift a payload of 1360 kg. The gas to be used is Helium at

101.3 Kpa and 23 0 C. The surrounding air is 101.3 KPa and 10 0 C.


1
Expert's answer
2021-11-29T11:42:09-0500

"F_A=F,"

"\rho g V=mg,"

"\rho V=m,"

"V=\frac m{\rho},"

"V=\frac{\pi d^3}6,"

"d=\sqrt[3]{\frac{6m}{\pi \rho}},"

"d=24.7~m."